An example that uses the elimination method to solve two linear equations.
Please help this is what I got. Not sure if it is right.
x + 6 = 11
–6 –6
x = 5

Answer by Alan3354(69443)

x + 6 = 11
–6 –6
x = 5
---------
To check it, sub 5 for x in the original eqn:
x + 6 = 11
5 + 6 = 11
It's right.